OpenStudy (anonymous):

Medal****** y - 1= (4y - 2)/3

OpenStudy (anonymous):

@undeadknight26

OpenStudy (anonymous):

@One098

OpenStudy (texaschic101):

y -`1 = (4y - 2) / 3 --- multiply both sides by 3 3(y - 1) = 4y - 2 can you see how I got that...the 3's on the right side cancel out. Now we distribute through the parenthesis 3y - 3 = 4y - 2 -- subtract 4y 3y - 4y - 3 = -2 -- add 3 3y - 4y = -2 + 3 -- combine terms -y = 1 -- multiply by -1 to make y positive y = -1 check.. y - 1 = (4y - 2) / 3 -1 - 1 = (4(-1) - 2) / 3 -2 = (-4 - 2) / 3 -2 = -6/3 -2 = -2 (correct) So y = -1 any questions ?